MODS1 AGw Commissioning at LDG: 16 November 2010 UTC

Observer(s): R. Pogge and B. Shappee (OSU), J.Hill and M.Pedani (LBTO)
Telescope Operator: S. Allanson (LBTO)
Active Optics Support: D. Miller (LBTO, remote from Tucson)
Telescope Support: M. Midkiff (LBTO)
Instrument Support: T. Shih (LBTO)

Summary

Relatively clear with some haze in the late afternoon, some puffy clouds around coming in from the NW (the system is off to the NE of us). Cooler than yesterday (T=-2C), but more humidity (66%), and wind 10 m/sec with gusts to 16 m/sec at 4pm in the afternoon.

Measured the LEFTZEROPOINT correctly and changed the PCS config. Verified it on a number of pointing and offset tests. PErformed initial tests of slit acquisition in RADEC and DETXY offsets to verify the system is performing correctly with no overt sign errors in the basic transformations. Successfullly acquired and guided while taking a 10-minute spectral observation, so the system is stable at least on timescales of typical long-dwell integrations.

Measured standard stars with the guider and WFS cameras (the latter not so successful as seeing degraded), to evaluate the photometric performance of both.

With the correct rotator zero point, shot a Stone, Pier, and Monet astrometric calibration field to take a first cut at the blue and red channel astrometric mapping.

Tested the transformation data acquisition script during seeing too poor for a formal data gathering run.

Built, installed, and tested a new collimation table for MODS1 at LDG

Preparations

LN2 fill at 2240 UTC, still liquid in the tanks, so hold times are >24h

James Howard came up during the day to check the HBS system, our problems last night were pressure and cooling issues. James increased pressure, changed filters, and worked on a chattering bypass valve. We should be OK for tonight.

Transformation measurements will be done in the SDSS r filter, so the reference wavelength for the MODS images is 0.6-microns. The effective guider wavelength is 0.575-microns (difference from 0.6-micron is not that important).

Details

Startup

0020 - Nearing sunset, doors open. T=-3C, RH=76% Windw 10-12m/sec, gusts to 17m/s

A few thin hazy clouds on the far western horizon, otherwise looking to be a nice night, if blustery. We'll have to watch the wind. It is colder than yesterday, and the mirrors are still cold from last night, so we should catch up fast.

0030 - First preset at twilight for pointing check

Coefficients:
   IE=+2
   CA=+35

PCS parameters updated
  LEFTPOINTINGMODEL = Sept26 model
  LEFTWAVELENGTH = 0.622 

Collimation

0050 - Pointed to BS9133, on-axis active preset to set collimation.

Preset failed, didn't acquire star. Problem was IE and CA were not re-entered after a PCS restart to enter parameters for MODS.

0100 - Re-ran the preset, much better.

Acquired, starting corrections. One wrinkle was that we started in split mode, changed to primary-only mode for corrections.

0107 - resent the preset after changing to primary only.

0112 - Hit the Y-limit on the primary. Taking off 1mm and resending the preset

Converged, getting 1.1-1.2 arcsec seeing. Temperature of the mirrors and structure are better behaved than last night.

0130 - Wait for twilight, watching the wind, which is having some high gusts flirting with the 20m/s limit.

Rotator Zero Point Measurement

0135 - Current value of LEFTZEROPOINT in PCSInstrument.conf is 30.157 for MODS1, which is the value found during September

Because of the ISS mounting issues, we are remeasuring this.

0137 - Image of the LS5x60x0.6 long slit mask with IMCS correction, dual mode, 60s, SDSS g and r filters.

Image: mods1x.20101116.0001 (will omit the prefixes except when necessary hereafter).

0140 - Second image. First image had forgotten to recenter the IMCS after a long dwell on-sky, this one is done right.

Of course, I really should have done a preset with the guide probe off-axis, so these images are useless. Some startup goofs when trying to interlace observing and instruction for the incoming student...

0143 - Stone O field, PA=0, guide star "1136", sent as active preset

Will execute script grid25.pro, runs a 5x5 grid with 20-arcsec spacing with absolute offsets.

Seeing on GCS 1.0-1.2 arcsec

0150 - Started the 5x5 grid. Seeing around 1arcsec, but variable, and wind is still gusty (15m/s, gusts to 19m/s)

Images: 0003 through 0027

Oops - left the slit mask in for the first image, will repeat at the end since the script returns to (0,0,0) absolute. This will become our slit image (having blown images 0001 and 0002).

Executing just fine, seeing is variable, 0.9 at best, 1.1-1.2 typical.

0225 - Seeing beginning to degrade, now about 2arcsec or worse! Seeing it in the grid images. Humidity approaching 80% wind still 15m/sec, gusts up to 18m/s

0227 - Sequence done, reshooting 0,0 (will be image 0028 in both channels)

0230 - Reshooting the slit image, 60s in both channels, image 0029 in both channels.

Problem: We realized after taking the data that we had left the last preset in active mode, rather than track mode, so we can't use these data for the LEFTZEROPOINT (although with some massaging it gives us the guide probe XY stage rotation). Will repeat later.

Guide Probe Photometric Sensitivity

0240 - Humidity rising relatively fast, and seeing degrading, now 2.5 arcsec FWHM. Getting warnings as we get near 90%. Watching...

0320 - IC1613, subfield A and B (our finder notation)

Procedure
  • Active preset on a bright star, tweak up collimation, stop guiding, capture a full-frame GCam image
  • Acquire preset on fainter stars (seeing too poor for guide lock), stop guiding, capture a GCam image

Repeated for the various photometric stars, clear filter on the guide camera.

We measure the bright stars twice: once behind the WFS pickoff beamsplitter, once on the center of the guide camera away from the beam splitter. This gives us a measurement of the attenuation of the beam splitter.

Seeing has been 1.6-2 arcsec, not so nice. Wind steady at 12m/sec with gusts to 18m/s, humidity hovering around 85%

Star # Pres.Mode left_guiderXXX.fits Filter Exp.Time (s)
1512 ACT 01 CL 1
1512 ACQ 02 CL 1
2048 ACQ 03 CL 10
2034 ACQ 04 CL 10
2019 ACQ 05 CL 10
1841 ACQ 06 CL 10
1917 ACQ 07 CL 10
2076 ACQ 08 CL 15
1512 ACT 09 CL 1
1512 ACQ 10 CL 1
1864 ACQ 11 CL 5
2107 ACQ 12 CL 15
1881 ACQ 13 CL 10
1854 ACQ 14 CL 10
2003 ACQ 15 CL 15
1887 ACQ 16 CL 5
1989 ACQ 17 CL 10
1843 ACQ 18 CL 10
1904 ACQ 19 CL 15
1512 ACT 20 B 15
1512 ACQ 21 B 5
1495 ACT 22 B 5
1495 ACQ 23 B 5
1831 ACQ 24 CL 10
1755 ACQ 25 B 30
1755 ACQ 26 LP 20
1755 ACQ 27 LP 25
1755 ACQ 28 LP 30
1755 ACQ 29 LP 35
1755 ACQ 30 LP 45
1755 ACQ 31 LP 50
1755 ACQ 32 LP 55
1755 ACQ 33 LP 65
1495 ACQ 34 LP 10
1495 ACQ 35 LP 20
1831 ACQ 36 LP 30
1932 ACQ 37 LP 10
2066 ACQ 38 LP 30
1999 ACQ 39 LP 30
FLAT ACQ 42 CL 30
FLAT ACQ 43 CL 30
FLAT ACQ 44 CL 30
FLAT ACQ 45 CL 30
FLAT ACQ 46 CL 30
FLAT ACQ 48 B 30
FLAT ACQ 49 B 30
FLAT ACQ 50 B 30
FLAT ACQ 51 B 30
FLAT ACQ 52 B 30
FLAT ACQ 53 LP 30
FLAT ACQ 54 LP 30
FLAT ACQ 55 LP 30
FLAT ACQ 56 LP 30
FLAT ACQ 57 LP 30

0425 - Completed observations. Images in : /OldRepository/20101116_MODS/GUIDE-PHOT
   Clear Filter - 16 stars
   F525LP Filter - 9 stars
   B_Bessel Filter - 3 stars (only brightest)

0436 - Sky flats dithered 5 (+ pattern) in each filter, 30s per frame. We are dithering by offsetting the guide probe by hand.

Conditions are still poor, seeing (instantaneous) 1.3 arcsec but variable. Humidity falling back towards 70%, Wind still 17-18m/s. No strong gusts above the limits for a long while.

Slit Scan - LEFTZEROPOINT Measurement

0500 - Active preset to Stone A2 with the "1507" guide star, tweaking up collimation. Seeing is about 1.7 arcsec but variable.

Then re-issuing the preset in track mode and running the slitscan.pro script +/-30-arcsec N/S offset in 15-arcsec intervals.

0505 - started sequence. Starts with mods1r.20101116.0030 through 0034 (no blue channel for this).

0512 - 0.6arcsec long-slit mask image 120s, image 0035

0525 - Weather update: T_air = -3.7, T_mirr = -2.9, RH ~ 69%, Wind ~ 19m/s, Seeing ~ 2.0-2.6"

0530 - Measured a rotation of the line of N/S star images relative to the Red CCD xy system is -2.25 degrees (cw), and the slit is rotated 0.28-degrees (ccw) so we need to rotate the instrument by 2.53-degrees clockwise to line the stars up with the slit.

The old LEFTZEROPOINT is 30.157 degrees. New LEFTZEROPOINT is 32.687 degrees. Did we get the sign right?

0535 - confirmatory slit scan

Reset the pointing first:
   IE = +29
   CA = +45

Acquiring images mods1r.20101116.0036-0040 with script slitscan.pro. Seeing is awful (3-arcsec).

Nope, wrong sign. This is why I don't gamble...

0544 - Setting LEFTZEROPOINT=27.63 degrees and running again

On doing an active preset to recover the field, they messed with the mirror C00, and we lost the star (massive defocus). Reset the preset.

Images: 0041-0045 (red channel)

0600 - Yep, that's the right answer. An overlaid slit image shows nice alignment.
   LEFTZEROPOINT=27.63

0608 - Running slitscan again, but now with an active preset and having changed the rotator zero point

Images: 0046-0050. Seeing about 1.5 arcsec during the run.

Looks very good. Any residual misalignment is due to the AGw-to-PCS_xy transformation (rotation of the AGw stage relative to the slit).

WFS Photometric Sensitivity

0628 - On-axis active presets to stars of known brightness in the Stone A2 field (because we're there).

Procedure:
  • active preset to a star on-axis, allow collimation to converge
  • stop guiding and wfs
  • acquires full-frame WFS images using readWFSCam

repeat for a range of brightnesses. Exposure times are chosen to get unsaturated Shack spots.

0708 - Seeing Crazy now! up to 5.5"

Star# Mode where Exp.Time left_wfscXX.fits
1183 ACT on-axis 30s 0002
1076 ACT on-axis 30s 0003
1076 ACT on-axis 10s 0004
1442 ACT on-axis 30s 0005
1265 ACT on-axis 10s 0006
1377 ACT on-axis 20s 0007
1545 ACT on-axis 50s 0008
1647 ACT on-axis 150s 0009

Data are here: /OldRepository/20101116_MODS/WFS-PHOT

07:14 - Done but data could be bad due to poor seeing

0730 - We're down to what we can do when the seeing is bad. We're at about midnight, and good news is that humidity is no longer a danger, and the wind shows some signs of trending to lower speed for the past hour. Seeing is about 1.5-2 arcsec but variable. Temperature has been dropping, though, down to -4.9, mirror and structure are now warmer than the outside air. Just not catching a break.

So, we talked about maybe doing a pointing map, but the telescope is out of thermal equilibrium, so it hardly seems useful

Spectral Target Acquisition - First Attempt

0800 - Doing a first cut at target acquisition with the long slit mask and grating mode, now that we have a good rotator zero point. Seeing is still poor, so we'll use a wide slit.

Prototype Acquisition Procedure:
  • Send active preset to target
  • during preset, insert the slit mask
  • On source - center up IMCS while the collimator is settling
  • Take an image through the slit mask with the red camera
  • Extract the slitmask during readout
  • Take a direct image of the target field.
  • Display the sum of slit+target
  • Compute the offset to move the star onto the slit
  • Execute the offset, putting the slit mask back in during the offset
  • Take an image through the mask - refine the centering as required
  • Put red channel into imaging mode, settle IMCS, start integrating.

We'll do this by hand to see how each step works and watch for hitches and bottlenecks.

Star: G191B2B, using the 1.2-arcsec slit because seeing is 1.5-2.0 arcsec FWHM and bouncy

Worked OK with RA/Dec offsets for PA=0

0830 - Trying a rotated acquisition, target is a faint object (fainter than a standard star)

Target was 3C120 (we had a finder), but at PA=-90, so targetting offsets had to be det_xy coordinates. This worked fine and acquisition was reasonably efficient considerintg it was our first attempt at it. Bottom line - no surprises with signs, scale, or other problems. We had a target 1.2" from the slit, issued the offset, and it popped in. Same as with the RADEC coordinate offsets we used with the PA=0 target above.

We took a couple of long spectra (120s and 600s, respectively), to see how well the system was tracking, the IMCS holding lock, etc. It behaved beautifully despite the poor seeing, which began to show signs of improvement during the sequence.

0933 - Wind still up (trended higher after a slight decline), humidity is tracking downward, and temperature as flattened out, so the telescope is finally starting to equilibrate again.

AGw Stage Transformation data collection

0950 - running a copy of Doug Miller's script for collecting AGw stage transformation data. We are in the Stone E field. John is running the IDL script. transform_collect, /MODS
Data is in /OldRepository/AGW_Data/20101116/ (different directory than that used above).

Just a quick test to verify we can run it. We still need to revisit the image headers and make sure Doug gets what he needs.

MODS Collimation Table

1005 - Running the stars in catalog Collimation-07hrs-Nov10.cat

1015 - Checked pointing as we were kind of far from the center on the last preset
   IE = 18
   CA = 32

1018 - Resuming the collimation table, the first start at dec=31 was too close to the zenith,

Short FITS images with the information in the headers are mods1r.20101116.0074 thru 101

Seeing was OK at the start, even got to 0.8 briefly, but on the way back up to high elevation it started to degrade. recovering at higher elevation. This night is a real teaser...

1124 - 28 images acquired by the end of the run, down and up in elevation.

Astrometric Field - Science Camera Astrometric Calibration

1128 - Stone E field, seeing variable 0.8 to 1.3

Images: mods1b.xxx.0037, mods1r.xxx.0103, g and r, 30sec, IMCS centered and locked on.

Run for PA=0, images mods1r.xxx.0102-0108 and mods1b.xxx.0036-0040
   g - 2x30s
   r - 2x30s
   i - 2x30s
   z - 2x30s
   u - 2x60s
Pairs: (g+r)x2, u+(ix2), u+(zx2)

1140 - Run for PA=180, same sequences, images mods1r.xxx.0109-... and mods1b.xxx.0041-...

New Collimation Table

1200 - Marco has a new collimation table based on the data taken above. We're taking it for a spin at the end of the night (since seeing is still too bad to do anything else useful).

1204 - We test the New Collimation Table

El=65 well rounded stars with ~1.1" seeing
EL=34 well rounded stars with ~1.2-1.4" seeing
EL=75 well rounded stars with ~1.0" seeing

End of Night Round Up

1217 - End of the night for us, we're done. Lots done, despite some trying conditions, and no big problems with the equipment.

1220 - Took bias images in the Guide and WFS cameras - dark instrument and 100msec exposures
    GCam bias images: 60-79
    WFS bias images: 13-32

1230 - That's all for the night.

-- RichardPogge - 15 Nov 2010
Topic revision: r17 - 16 Aug 2011, RichardPogge
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback